OpenStudy (anonymous):

How many ways can a jury of 6 women and 6 men be selected from 10 women and 12 men?

OpenStudy (anonymous):

\[\dbinom{10}{6}\times \dbinom{12}{6}\]

OpenStudy (anonymous):

or \[210\times 924\]

OpenStudy (anonymous):

194040

OpenStudy (anonymous):

so is it both the right answer or its one or the other..cuz isnt the fraction 120/6... and 210x924=194040

OpenStudy (anonymous):

??

OpenStudy (anonymous):

\[\dbinom{10}{6}=\frac{10\times 9\times 8\times 7}{4\times 3\times 2}\] the number of ways to choose 6 out of 10. that is 210

OpenStudy (anonymous):

\[\dbinom{12}{6}\] is the number of ways to choose 6 from 12. you may have seen this written as 12C6. that is 924 and their product is 194040